. the inflation rate[3].

6. UK’s real GDP was 1 360 trillion pounds [£] in 2009 and 1 434 trillion pounds in 2010. UK’s population was 191.5 million people in 2009 and 193.3 million in 2010.
Calculate.
a. the economic growth rate.[3]
b. The growth rate of real GDP per person [5]
c. The approximate number of years it takes for real GDP per person in the UK to double if the 2010 economic growth rate and population growth rate are maintained.[4]
1
Expert's answer
2019-08-14T09:26:42-0400

a."\frac{1434-1360}{1360}"=0.05441


b. "\frac{0.05441}{193.3-191.5}" = 0.03023


c. "\frac{72}{3}" =24 years